# HoH: A Dynamic Benchmark for Evaluating the Impact of Outdated Information on Retrieval-Augmented Generation

### v1

@@ Abstract
While Retrieval-Augmented Generation (RAG) has emerged as an effective approach for addressing the knowledge outdating problem in Large Language Models (LLMs), it faces a critical challenge: the prevalence of outdated information in knowledge bases. Current research primarily focuses on incorporating up-to-date information, yet the impact of outdated information coexisting in retrieval sources remains inadequately addressed. To bridge this gap, we introduce HoH, the first benchmark specifically designed to evaluate the impact of outdated information on RAG. Our benchmark leverages token-level diff algorithms combined with LLM pipelines to efficiently create a large-scale QA dataset that accurately captures temporal knowledge evolution in real-world facts. Through comprehensive experiments, we reveal that outdated information significantly degrades RAG performance in two critical ways: (1) it substantially reduces response accuracy by distracting models from correct information, and (2) it can mislead models into generating potentially harmful outputs, even when current information is available. Current RAG approaches struggle with both retrieval and generation aspects when handling outdated information. These findings highlight the urgent need for innovative solutions to address the temporal challenges in RAG.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing with their remarkable knowledge retention and reasoning abilities OpenAI et al. (2024); Grattafiori et al. (2024); Qwen (2024). However, the static nature of their pre-trained knowledge poses a significant challenge in handling rapidly evolving real-world information. LLMs frequently generate outdated responses that, while appearing plausible, no longer reflect current facts Maynez et al. (2020); Liu et al. (2023). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Guu et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022) has emerged as a promising solution to this knowledge outdating problem by dynamically retrieving and integrating up-to-date information from external sources.
Despite RAG’s promise, a critical challenge remains unaddressed: the pervasive presence of outdated information in knowledge bases Xin et al. (2024). As illustrated in Figure 1, when querying about the current US president, RAG systems may retrieve both current and outdated information, potentially leading to confusion or incorrect responses. This scenario is commonplace as information evolves and outdated facts inevitably accumulate across various sources, particularly in search engine scenarios where content is cached and redistributed. Current research primarily focuses on how to effectively retrieve and leverage the latest information Karpukhin et al. (2020); Chen et al. (2024), while overlooking how the presence of outdated information could harm RAG. This oversight is significant, as even state-of-the-art industry RAG systems Mehdi (2023) frequently generate incorrect responses due to outdated information Vu et al. (2024). To address this challenge, we introduce HoH  [^1] HoH resembles a shocked emoji [Uncaptioned image], expressing our alarm at the harm caused by outdated information. (How outdated information Harms Retrieval-Augmented Generation), the first large-scale benchmark designed to evaluate RAG’s robustness against outdated information.
HoH comprises two primary components: HoH-QA and HoH-SearchEngine. HoH-QA is a continuously evolving QA dataset that maintains temporal alignment with contemporary world knowledge. Our dataset features comprehensive evidence annotations for each QA pair, enabling fine-grained evaluation of individual RAG components. More importantly, we introduce annotations for outdated answers and evidence for the first time, allowing systematic assessment of how outdated information impacts RAG performance. Complementing the QA dataset, HoH-SearchEngine simulates real-world search scenarios by maintaining both current and historical documents, better reflecting the actual challenges RAG faced in practice. Our benchmark currently includes 96,124 QA pairs and 219,463 documents, far surpassing the scale of existing datasets.
@@
Table 1: Comparison of our HoH with existing dynamic benchmarks. The automation indicates the feasibility of automatic generation. The Maintenance represents whether the validity of previously generated datasets is verified in the forthcoming time step. The evidence text indicates whether the dataset includes the evidence text. The outdated info shows whether the dataset contains annotation for outdate information. The question number and article number identify the number of the dataset question and the number of the article for retrieval, respectively.
@@ 2.2 Present-Anchored Question Answering
As shown in Table 1, our work, HoH, builds upon these foundations while addressing several key limitations in existing benchmarks. Through automation and continuous maintenance, HoH ensures both scalability and data validity. HoH explicitly includes evidence text annotations, enabling fine-grained analyses of both retrieval and generation components in RAG systems. More importantly, it uniquely provides explicit annotations for outdated information, enabling systematic study of its impact on RAG. Lastly, HoH offers an unprecedented scale with 96,124 questions and 219,463 articles, substantially surpassing existing benchmarks in terms of data volume and comprehensiveness.
@@ 3.1 HoH-QA
The quality of generated QA pairs is ensured through a multi-stage review process. After initial generation, LLMs conduct a thorough review, focusing on the quality of both questions and answers. Detailed generation prompts and quality control procedures are provided in Appendix B.3.
@@ A.1 Effectiveness of Change Extraction
These results validate the effectiveness of our approach in identifying factual changes, which is fundamental to ensuring the quality of our HoH dataset.
@@ B.4 Details of Automatic Update
We omit how to distinguish between newly emerged sentence pairs and pre-existing sentence pairs in the main text. Taking sentence pair $S=(s_{old},s_{new})$ of article $A$ as an example. We first identify all QA pairs generated from article $A$ in the dataset and locate all corresponding evidence $E$. If the knowledge in the dataset has changed again, then there should be some evidence $e_{i}\in E$ that matches $s_{old}$. If any $e_{i}\in E$ matches $s_{old}$, we identify $S$ as a pre-existing sentence pair. For pre-existing sentence pairs, we subsequently employ prompt shown in Figure 8 to generate a new answer. The quality review prompt is the same as QA Generation.
@@ E.1 Vanilla Results
E.2 Additional Results for Generator
We conduct additional experiments on the effect of passage types on generation performance. As shown in Table 14, when only outdated passages($\mathcal{O}$) are available without relevant $\mathcal{R}$), the generator’s performance indicate that LLMs are inevitably misled by outdated information, producing a high proportion of potentially harmful responses.
Moreover, we conduct additional experiments focusing on long answers. As demonstrated in Table 15 and Table 16, the experimental findings largely align with those observed in short-answer scenarios. The impact of distracting passages($\mathcal{D}$) remains substantially lower than that of outdated passages($\mathcal{O}$), and all three models exhibit significant sensitivity to passage ordering. One notable difference is that Qwen-7B demonstrates marked improvement in long-answer performance, with its optimal performance even surpassing Llama-8B under appropriate passage sequencing conditions.

### v2

@@ Abstract
While Retrieval-Augmented Generation (RAG) has emerged as an effective approach for addressing the knowledge outdating problem in Large Language Models (LLMs), it still faces a critical challenge: the prevalence of outdated information in knowledge bases. Current research primarily focuses on incorporating up-to-date information, yet the impact of outdated information coexisting in retrieval sources remains inadequately addressed. To bridge this gap, we introduce HoH, the first benchmark specifically designed to evaluate the impact of outdated information on RAG. Our benchmark leverages token-level diff algorithms combined with LLM pipelines to efficiently create a large-scale QA dataset that accurately captures the evolution of temporal knowledge in real-world facts. Through comprehensive experiments, we reveal that outdated information significantly degrades RAG performance in two critical ways: (1) it substantially reduces response accuracy by distracting models from correct information, and (2) it can mislead models into generating potentially harmful outputs, even when current information is available. Current RAG approaches struggle with both retrieval and generation aspects when handling outdated information. These findings highlight the urgent need for innovative solutions to address the temporal challenges in RAG. Our code and data are available at https://github.com/0russwest0/HoH.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing with their remarkable knowledge retention and reasoning abilities OpenAI et al. (2024); Grattafiori et al. (2024); Qwen (2024). However, the static nature of their pre-trained knowledge poses a significant challenge in handling rapidly evolving real-world information. LLMs frequently generate outdated responses that, while appearing plausible, no longer reflect current facts Maynez et al. (2020); Liu et al. (2023). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Guu et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022); yu2024knowledge has emerged as a promising solution to this knowledge-outdating problem by dynamically retrieving and integrating up-to-date information from external sources.
Despite RAG’s promise, a critical challenge remains unaddressed: the pervasive presence of outdated information in knowledge bases Xin et al. (2024). As illustrated in Figure 1, when querying about the current US president, RAG systems cheng2025survey; ouyang2024revisiting may retrieve both current and outdated information, potentially leading to confusion or incorrect responses. This scenario is commonplace as information evolves and outdated facts inevitably accumulate across various sources, particularly in search engine scenarios where content is cached and redistributed. Current research primarily focuses on how to effectively retrieve and leverage the latest information Karpukhin et al. (2020); Chen et al. (2024), while overlooking how the presence of outdated information could harm RAG. This oversight is significant, as even state-of-the-art industry RAG systems Mehdi (2023) frequently generate incorrect responses due to outdated information Vu et al. (2024). To address this challenge, we introduce HoH  [^1] HoH resembles a shocked emoji [Uncaptioned image], expressing our alarm at the harm caused by outdated information. (How outdated information Harms Retrieval-Augmented Generation), the first large-scale benchmark designed to evaluate RAG’s robustness against outdated information.
HoH comprises two primary components: HoH-QA and HoH-SearchEngine. HoH-QA is a continuously evolving QA dataset that maintains temporal alignment with contemporary world knowledge. Our dataset features comprehensive evidence annotations for each QA pair, enabling fine-grained evaluation of individual RAG components. More importantly, we introduce annotations for outdated answers and evidence for the first time, allowing systematic assessment of how outdated information impacts RAG performance. Complementing the QA dataset, HoH-SearchEngine simulates real-world search scenarios by maintaining both current and historical documents, better reflecting the actual challenges RAG faces in practice. Our benchmark currently includes 96,124 QA pairs and 219,463 documents, far surpassing the scale of existing datasets.
@@
Table 1: Comparison of our HoH with existing dynamic benchmarks. The automation indicates the feasibility of automatic generation. The Maintenance represents whether the validity of previously generated datasets is verified in the forthcoming time step. The evidence text indicates whether the dataset includes the evidence text. The outdated info shows whether the dataset contains annotation for outdate information. The ’question number’ specifically counts Time-Sensitive Questions (excluding time-irrelevant ones), and ’article number’ identifies the number of articles for retrieval, respectively.
@@ 2.2 Present-Anchored Question Answering
As shown in Table 1, our work, HoH, builds upon these foundations while addressing key limitations in existing benchmarks. Through automation and continuous maintenance, HoH ensures both scalability and data validity. It includes evidence text annotations, enabling fine-grained analyses of both retrieval and generation components in RAG systems. More importantly, it uniquely provides annotations for outdated information, enabling systematic study of its impact. Lastly, HoH offers unprecedented scale with 96,124 questions and 219,463 articles, surpassing existing benchmarks in comprehensiveness.
@@ 3.1 HoH-QA
The quality of generated QA pairs is ensured through a multi-stage review process. After initial generation, LLMs conduct a thorough review, focusing on the quality of both questions and answers. Detailed generation prompts, quality control procedures, and a manual validation confirming the high reliability of this automated process are provided in Appendix B.3.
@@ 7 Limitation
Acknowledgments
This research was supported by grants from the National Key Research and Development Program of China (Grant No. 2024YFC3308200), the grants of the Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities (No. WK2150110032).
@@ A.1 Effectiveness of Change Extraction
These results validate the effectiveness of our approach in identifying factual changes, which is fundamental to ensuring the quality of the constructed HoH dataset.
@@ B.3 Details of QA Generation
To further validate the reliability of this LLM-based quality review process, we conducted an additional manual evaluation. We randomly selected 500 QA pairs from the HoH-QA dataset for this manual assessment. Each QA pair was evaluated by human annotators based on the four primary quality dimensions covered by our automated pipeline: Same Answer Check, Answer Accuracy, Question Clarity and Completeness, and Temporal Independence. The results of this manual evaluation, presented in Table 9, demonstrate a high level of reliability in our automated quality assurance procedures.
Table 9: Manual evaluation results for the LLM-based quality review of 500 QA pairs, assessing key quality dimensions.
Evaluation Perspective Accuracy (%)
Same Answer Check 100.0
Answer Accuracy 99.2
Question Clarity and Completeness 98.4
Temporal Independence 100.0
@@ B.4 Details of Automatic Update
We omit how to distinguish between newly emerged sentence pairs and pre-existing sentence pairs in the main text. Taking sentence pair $S=(s_{old},s_{new})$ of article $A$ as an example. We first identify all QA pairs generated from article $A$ in the dataset and locate all corresponding evidence $E$. If the knowledge in the dataset has changed again, then there should be some evidence $e_{i}\in E$ that matches $s_{old}$. If any $e_{i}\in E$ matches $s_{old}$, we identify $S$ as a pre-existing sentence pair. For pre-existing sentence pairs, we subsequently employ the prompt shown in Figure 8 to generate a new answer. The quality review prompt is the same as QA Generation.
@@ E.1 Vanilla Results
E.2 Additional Results for Retrieval
We conduct additional experiments to evaluate existing temporal retrieval solutions, MRAGsiyue2025mragmodularretrievalframework and TempRALMgade2024itstimeincorporatingtemporality, comparing them against our baseline retrieval approach. These methods aim to improve retrieval by incorporating temporal factors. As shown in Table 15, while MRAG effectively filters outdated passages ($\mathcal{O}$), it significantly lowers the hit rate of relevant passages ($\mathcal{R}$). TempRALM offers a better balance, but the reduction in $\mathcal{R}$ hit rate remains substantial. These results highlight the inherent conflict between maximizing relevant information retrieval and minimizing outdated content retrieval, a challenge that current temporal-focused methods do not fully resolve without notable sacrifices in relevant content recall.
Table 15: Hit rate of relevant ($\mathcal{R}$) and outdated ($\mathcal{O}$) passages for the baseline retrieval method compared to temporal retrieval methods.
Embedding Type Type Baseline MRAG TempRALM
@5 @10 @5 @10 @5 @10
bge-base-en-v1.5 $\mathcal{R}$ 0.7019 0.7365 0.2921 0.3213 0.4473 0.4906
$\mathcal{O}$ 0.5289 0.5578 0.0534 0.0872 0.2264 0.2613
bge-m3 $\mathcal{R}$ 0.7312 0.7578 0.3183 0.3464 0.4733 0.5143
$\mathcal{O}$ 0.5507 0.5698 0.0574 0.0989 0.2399 0.2718
E.3 Additional Results for Generation
We conduct additional experiments on the effect of passage types on generation performance using the open-source models detailed in our main experiments. As shown in Table 18 (short-answer scenarios), when only outdated passages($\mathcal{O}$) are available without relevant $\mathcal{R}$), the generator’s performance indicates that these LLMs are inevitably misled by outdated information, producing a high proportion of potentially harmful responses.
Moreover, we conduct further experiments with these open-source models focusing on long answers. As demonstrated in Table 19 and Table 20, the experimental findings largely align with those observed in short-answer scenarios. The impact of distracting passages($\mathcal{D}$) remains substantially lower than that of outdated passages($\mathcal{O}$), and all three tested open-source models (e.g., Llama 3.1 70B, Llama 3.1 8B, Qwen 2.5 7B – *adjust model names if these are not the ones in those specific tables*) exhibit significant sensitivity to passage ordering. One notable difference is that Qwen-7B demonstrates marked improvement in long-answer performance, with its optimal performance even surpassing Llama-8B under appropriate passage sequencing conditions.
Building on these observations with open-source models, and to broaden the scope of our generator analysis, we also evaluated a leading proprietary model, GPT-4o. While our primary experiments centered on open-source LLMs due to budget considerations, these supplementary experiments with GPT-4o, conducted on a representative subset of 1,000 test samples, provide insights into how these challenges manifest in highly capable closed-source systems. The results for GPT-4o concerning distracting passages are presented in Table 16, and those detailing the impact of outdated information and passage ordering are in Table 17.
The findings for GPT-4o are entirely consistent with the main conclusions drawn in our paper regarding generator vulnerabilities. GPT-4o is also highly sensitive to outdated information ($\mathcal{O}$), exhibiting significantly greater performance degradation when exposed to $\mathcal{O}$ compared to distracting passages ($\mathcal{D}$). Overall, GPT-4o’s performance is comparable to that of Llama 3.1 70B. It performs slightly better than Llama 3.1 70B in scenarios without outdated passages (as seen in Table 16), but slightly worse when outdated passages are present (Table 17). Furthermore, passage ordering strategies also have a substantial impact on GPT-4o’s performance, although the magnitude of this effect is somewhat less pronounced than for Llama 3.1 70B.
Table 16: Performance of GPT-4o with passages sorted by relevance score in descending order for {×$\{\mathcal{R}\times 1,\mathcal{D}\times n\}$. Per. (Perfect), Mis. (Missing), Har. (Harmful) and Score are in percentage.
$n$ Perfect Missing Harmful Score
1 93.6 2.3 4.1 89.5
2 92.8 2.8 4.4 88.4
3 93.0 2.1 4.9 88.1
4 93.0 2.4 4.6 88.4
6 93.1 1.8 5.1 88.0
[…]
Order $n$ Perfect Missing Harmful Score
[…]
1 78.4 6.5 15.1 63.3
2 78.7 5.4 15.9 62.8
3 79.0 5.0 16.0 63.0
5 79.9 3.8 16.3 63.6
[…]
1 78.8 5.4 15.8 63.0
2 79.7 3.5 16.8 62.9
3 78.3 3.7 18.0 60.3
5 78.8 3.3 17.9 60.9
[…]
1 80.3 7.7 12.0 68.3
2 80.3 6.2 13.5 66.8
3 79.9 5.9 14.2 65.7
5 78.2 5.0 16.8 61.4